OpenStudy (anonymous):

please help me out. :) what is x^2+y^2-2x-2y+1=0 in polar form?

OpenStudy (ksaimouli):

okay, so \[x=rcos \theta, y=rsin \theta,\]

OpenStudy (anonymous):

yes.. :)

OpenStudy (ksaimouli):

\[(rcos \theta)^2+(rsin \theta)^2-2(rcos \theta)-2(rsin \theta)+1=0\]

OpenStudy (anonymous):

okay, then what's next?

OpenStudy (ksaimouli):